OpenStudy (madmerc):

Solve x^2 + 4x + 8 = 0.

OpenStudy (anonymous):

you mind getting a complex number as an answer?

OpenStudy (dayakar):

do u have any idea about solution

OpenStudy (anonymous):

use the quadratic formula ... if anything troubles you tell us

OpenStudy (anonymous):

you can find the quadratic formula anywhere online

OpenStudy (dayakar):

(x^2 +2*x*2 + 2^2 ) +4 =0 [making the given equation in (a+b)^2 form] (x+2)^2 +4=0

OpenStudy (anonymous):

i wouldn't 4 is even so i would complete the square

OpenStudy (madmerc):

the answer choices are all in x + number plus/minus form so

OpenStudy (dayakar):

do u understand what i written

OpenStudy (madmerc):

@dayakar no i dont

OpenStudy (dayakar):

do u know (a+b)^2 formula

OpenStudy (madmerc):

not really @dayakar

OpenStudy (dayakar):

(a+b)^2 = a^2+2ab+ b^2

OpenStudy (dayakar):

is it clear

OpenStudy (madmerc):

no I dont understand what that means @dayakar

OpenStudy (madmerc):

@satellite73 can you help

OpenStudy (madmerc):

okay the answer is \[x = -2 \pm 2i\]